OpenStudy (sugardemonhunters):

Why is my answer not correct? I think its 1/2 but the answer key says the answer is 2/45. Can someone explain why? Thank you :) A bag of fruit contains 3 apples, 2 oranges, 1 banana, and 4 pears. Gerald will randomly select two pieces of fruit one at a timefrom the bag and not put them back. What is the probability that the first piece of fruit gerald selects will be a banana and the second piece of fruit will be a pear

OpenStudy (sugardemonhunters):

I got 1/2 by adding 1(banana) and 4(pear) which would be 5/10 (total amount of fruits) which is equivalent to 1/2

zepdrix (zepdrix):

Calculate the odds of getting a banana, then separately, the odds of getting a pear after that, and multiply the results together.

zepdrix (zepdrix):

Banana = 1/10 Pear = 4/9 (Because we've removed one of the 10 pieces of fruit already)

OpenStudy (faiqraees):

No thats the wrong method. To calculate the probabilty of two "related event" you have to multiply the probabiltiy of getting a banana by the probability of getting a pear

OpenStudy (faiqraees):

Now the probability of getting a banana = 1/10 (since 1 banana and 10 fruits) probability of getting a pear = 4/9 (since 4 pears and 9 fruits ) Probability = 4/9 x 1/10

OpenStudy (sugardemonhunters):

Thank you all !